which of the following is the equation of the line that has tha same slope as y=-3/2x +2 and the same y-intercept as y=3x-2?
a)y-2=-3/2x
b)-3/2 x=y+2
c)y+2=-3/2
D)-3/2x=y+3

Answer by stanbon(75887)

which of the following is the equation of the line that has tha same slope as y=-3/2x +2 and the same y-intercept as y=3x-2?
a)y-2=-3/2x
y = (-3/2)x+2
No.
-----------------------
b)-3/2 x=y+2
y = (-3/2)x-2
No.
---------------
c)y+2=-3/2
y = -3/2-2 = -7/2
No
---------------
D)-3/2x=y+3
y = (-3/2)x-3
No
========================
Cheers,
Stan H.
================